# Fast Text-to-3D-Aware Face Generation and Manipulation via Direct Cross-modal Mapping and Geometric Regularization

### v1

@@ Abstract
Text-to-3D-aware face (T3D Face) generation and manipulation is an emerging research hot spot in machine learning, which still suffers from low efficiency and poor quality. In this paper, we propose an End-to-End Efficient and Effective network for fast and accurate T3D face generation and manipulation, termed $E^{3}$-FaceNet. Different from existing complex generation paradigms, $E^{3}$-FaceNet resorts to a direct mapping from text instructions to 3D-aware visual space. We introduce a novel Style Code Enhancer to enhance cross-modal semantic alignment, alongside an innovative Geometric Regularization objective to maintain consistency across multi-view generations. Extensive experiments on three benchmark datasets demonstrate that $E^{3}$-FaceNet can not only achieve picture-like 3D face generation and manipulation, but also improve inference speed by orders of magnitudes. For instance, compared with Latent3D, $E^{3}$-FaceNet speeds up the five-view generations by almost 470 times, while still exceeding in generation quality. Our code are released at https://github.com/Aria-Zhangjl/E3-FaceNet.
@@ 1 Introduction
Text-to-3D-aware face (T3D Face) generation is an emerging research hot spot in machine learning (Toshpulatov et al., 2021; Zhang et al., 2023b; Wu et al., 2023; Xia & Xue, 2023). Compared with the well-developed 2D face generation (Xia et al., 2021; Patashnik et al., 2021; Peng et al., 2022a, b), T3D Face not only requires to depict the intricate details of different facial components but also to model the 3D shapes and positions under different poses (Toshpulatov et al., 2021; Xia & Xue, 2023). Additionally, the semantic consistency between cross-modal and cross-view generations is also a main obstacle (Wu et al., 2023).
@@ 2 Related Work
The task of 3D-aware image generation (Xia & Xue, 2023) is to generate high-quality renderings that are consistent across multiple views. Recent studies have been conducted in representing 3D scenes with neural implicit functions(Mescheder et al., 2019; Park et al., 2019; Mildenhall et al., 2021), among which neural implicit representation (NIR) (Mescheder et al., 2019; Mildenhall et al., 2021; Park et al., 2019) emerges as an effective alternative for learning-based 3D reconstruction with multi-view 2D supervision. This representation is then applied to 3D image synthesis under single-view image supervision (Schwarz et al., 2020; Chan et al., 2021; Gu et al., 2021; Or-El et al., 2022; Chan et al., 2022). In this paper, we extend the representative implicit method, i.e., StyleNeRF (Gu et al., 2021), for text-conditioned 3D face generation. By incorporating text semantics into the latent code, our proposed methods enable the end-to-end mapping from text words to 3D faces.
@@ 2.2 Text-to-3D Face Generation
Text-to-3D face (T3D Face) generation often aims to incorporate text information into 3D generative models for language-guided 3D face generation. Dreamface (Zhang et al., 2023b) and Describe3D (Wu et al., 2023) propose to generate text-conditioned texture maps to render 3D morphable models (3DMM) (Blanz & Vetter, 2023). Despite the effectiveness of geometric modeling, these methods require additional processes to add more facial assets like hair or wrinkle. A relevant work to this paper is TG-3DFace (Yu et al., 2023), which uses tri-plane neural representations and extends EG3D (Chan et al., 2022), a 3D-aware GAN, for end-to-end text-conditioned generation. However, its text-control mechanism is still compositional, which needs to additionally integrate a face parser and train an attribute classifier for fine-grained semantics injection. To guide the generator in learning correct priors, TG-3DFace incorporates a camera pose conditional discriminator during the training process. Different from previous methods, our proposed $E^{3}$-FaceNet seamlessly integrated text information into the whole process of 3D representation and 2D upsampling, achieving single-stage cross-modal mapping without any further interventions. Besides, we also propose a set of innovative designs to cope with the problems encountered in direct T3D Face generations, i.e., 3D geometric regularization and Style Code Enhancer.
@@ 3.1 Preliminary
where $\phi_{\boldsymbol{w}}^{n}(\mathbf{x})$ is the $n^{th}$ feature of point $\mathbf{x}$, $f$ is the mapping network that projects the noise vector $z$ to the style space $\mathcal{W}$, and $g_{w}^{i}(.)$ is the $i^{th}$ MLP modulated by $w$.
@@ 3.3 Style Code Enhancer
As discussed above, it is often difficult to control the entire generation with a single style code $w$ in 2D image generation (Abdal et al., 2020; Saha et al., 2021). And this problem will become more prominent in T3D face generation, since the 3D visual semantic space is much larger than that of the 2D one (Xia & Xue, 2023). In this case, we introduce a Style Code Enhancer (SCE) to improve the impact of style codes in the 3D rendering process.
@@
Equipped with Style Code Enhancer, $E^{3}$-FaceNet can better embed fine-grained text information into the rendering process, while maintaining the image quality.
@@ 4.2 Experimental Results
Comparison with T3D Face methods. We first quantitatively compare our $E^{3}$-FaceNet with two strong T3D Face methods, namely Latent3D (Canfes et al., 2023) and Describe3D (Wu et al., 2023) on CelebAText (Sun et al., 2021) and FFHQ-Text (Zhou, 2021), of which results are given in Tab.1. From this table, we can first find that $E^{3}$-FaceNet greatly outperforms the other two methods under the metric of MVIC, e.g., +6.87% on CelebAText, showing great advantages in terms of multi-view identity preservation. Meanwhile, the performance of SA is also outstanding, which is slightly worse than Latent3D on CelebA but obviously better on FFHQ. These results suggest that $E^{3}$-FaceNet can achieve excellent multi-view semantic alignment. Moreover, retaining high performance, our $E^{3}$-FaceNet also speeds up the five-view generations by orders of magnitudes. For example, compared with Latent3D and Describe3D, the inference speed of $E^{3}$-FaceNet is 471.75 $\times$ and 35.83 $\times$ faster. Notably, $E^{3}$-FaceNet achieves such performance under the setting of zero-shot generation.
@@
Text-Driven 3D Face Manipulation. To examine the editing ability of $E^{3}$-FaceNet, we also compare it with Latent3D and ClipFace (Aneja et al., 2023) in Tab.2. Without available evaluations, we conduct a user study for this comparison. We can first observe that the identity preservation and the editing quality of $E^{3}$-FaceNet are much superior to the compared methods, e.g., 69.93 v.s. 21.92 of Latent3D on IP and 75.92 v.s. 12.48 of ClipFace on EQ. Besides, $E^{3}$-FaceNet also exhibits strong cross-modal semantic consistency, e.g., 52.64 v.s. 44.56 of ClipFace on SA. In addition, the efficiency of $E^{3}$-FaceNet is still much better than the others. For instance, ClipFace requires about 17 mins to train the texture mappers with another 1.8 seconds for rendering. In contrast, $E^{3}$-FaceNet only takes 0.89 seconds for the five-view editing.
@@ 5 Conclusion
Acknowledge
This work was supported by National Science and Technology Major Project (No. 2022ZD0118201), the National Science Fund for Distinguished Young Scholars (No.62025603), the National Natural Science Foundation of China (No. U21B2037, No. U22B2051, No. 62176222, No. 62176223, No. 62176226, No. 62072386, No. 62072387, No. 62072389, No. 62002305 and No. 62272401), and the Natural Science Foundation of Fujian Province of China (No.2021J01002, No.2022J06001).
@@ Appendix A Introduction of Comparison Methods
TediGAN (Xia et al., 2021) can perform text-guided 2D image generation by optimizing a random sampled noise $*z*$ via
@@ B.4.3 Evaluation Criteria
For each editing instruction, we render manipulated 3D faces from three different views: left, front, and right. Participants are presented with the original face and all editing results. We invite 25 individuals to assess the edited faces based on three criteria: (1) Identity Preservation (IP), selecting the edit result that best preserves the facial identity of the original face; (2) Semantic Alignment (SA), selecting the edit result that most closely aligns with the given instructions; and (3) Editing Quality (EQ), selecting the edit result with the highest perceptual quality. We calculated the percentage of times each method was chosen as the best performer by users, based on the previous mentioned criteria. As shown in Figure 8, the edit results are presented anonymously to the users.

### v3

@@ Abstract
Text-to-3D-aware face (T3D Face) generation and manipulation is an emerging research hot spot in machine learning, which still suffers from low efficiency and poor quality. In this paper, we propose an End-to-End Efficient and Effective network for fast and accurate T3D face generation and manipulation, termed $E^{3}$-FaceNet. Different from existing complex generation paradigms, $E^{3}$-FaceNet resorts to a direct mapping from text instructions to 3D-aware visual space. We introduce a novel Style Code Enhancer to enhance cross-modal semantic alignment, alongside an innovative Geometric Regularization objective to maintain consistency across multi-view generations. Extensive experiments on three benchmark datasets demonstrate that $E^{3}$-FaceNet can not only achieve picture-like 3D face generation and manipulation, but also improve inference speed by orders of magnitudes. For instance, compared with Latent3D, $E^{3}$-FaceNet speeds up the five-view generations by almost 470 times, while still exceeding in generation quality. Our code is released at https://github.com/Aria-Zhangjl/E3-FaceNet.
@@ 1 Introduction
Text-to-3D-aware face (T3D Face) generation is an emerging research hot spot in machine learning (Toshpulatov et al., 2021; Zhang et al., 2023b; Wu et al., 2023; Xia & Xue, 2023). Compared with the well-developed 2D face generation (Xia et al., 2021; Patashnik et al., 2021; Peng et al., 2022a, b; Du et al., 2023), T3D Face not only requires to depict the intricate details of different facial components but also to model the 3D shapes and positions under different poses (Toshpulatov et al., 2021; Xia & Xue, 2023). Additionally, the semantic consistency between cross-modal (Zhou et al., 2019; Luo et al., 2022; Zhou et al., 2021; Wu et al., 2023) and cross-view (Or-El et al., 2022; Chan et al., 2022; Xia & Xue, 2023) generations is also a main obstacle.
@@ 2 Related Work
The task of 3D-aware image generation (Xia & Xue, 2023) is to generate high-quality renderings that are consistent across multiple views. Recent studies have been conducted in representing 3D scenes with neural implicit functions(Mescheder et al., 2019; Park et al., 2019; Mildenhall et al., 2021), among which neural implicit representation (NIR) (Mescheder et al., 2019; Mildenhall et al., 2021; Park et al., 2019) emerges as an effective alternative for learning-based 3D reconstruction with multi-view 2D supervision. This representation is then applied to 3D image synthesis under single-view image supervision (Schwarz et al., 2020; Chan et al., 2021; Gu et al., 2021; Or-El et al., 2022; Chan et al., 2022). In this paper, we extend the representative implicit method, i.e., StyleNeRF (Gu et al., 2021), for text-conditioned 3D face generation. By incorporating text semantics into the latent code, our proposed method enables the end-to-end mapping from text words to 3D faces.
@@ 2.2 Text-to-3D Face Generation
Text-to-3D face (T3D Face) generation often aims to incorporate text information into 3D generative models for language-guided 3D face generation. Dreamface (Zhang et al., 2023b) and Describe3D (Wu et al., 2023) propose to generate text-conditioned texture maps to render 3D morphable models (3DMM) (Blanz & Vetter, 2023). Despite the effectiveness of geometric modeling, these methods require additional processes to add more facial assets like hair or wrinkle. A relevant work to this paper is TG-3DFace (Yu et al., 2023), which uses tri-plane neural representations and extends EG3D (Chan et al., 2022), a 3D-aware GAN, for end-to-end text-conditioned generation. However, its text-control mechanism is still compositional, which needs to additionally integrate a face parser and train an attribute classifier for fine-grained semantics injection. To guide the generator in learning correct priors, TG-3DFace incorporates a camera pose conditional discriminator during the training process. Different from previous methods, our proposed $E^{3}$-FaceNet seamlessly integrates text information into the whole process of 3D representation and 2D upsampling, achieving single-stage cross-modal mapping without any further interventions. Besides, we also propose a set of innovative designs to cope with the problems encountered in direct T3D Face generations, i.e., 3D geometric regularization and Style Code Enhancer.
@@ 3.1 Preliminary
where $\phi_{\boldsymbol{w}}^{n}(\mathbf{x})$ is the $n^{th}$ layer feature of point $\mathbf{x}$, $f$ is the mapping network that projects the noise vector $z$ to the style space $\mathcal{W}$, and $g_{w}^{i}(.)$ is the $i^{th}$ MLP modulated by $w$.
@@ 3.3 Style Code Enhancer
As discussed above, it is often difficult to control the entire generation with a single style code $w$ in 2D image generation (Abdal et al., 2020; Saha et al., 2021). And this problem will become more prominent in T3D face generation, since the 3D visual semantic space is much larger than that of the 2D one (Xia & Xue, 2023). Previous studies (Karras et al., 2019; Patashnik et al., 2021; Xia et al., 2022) have shown the varying levels of detail captured by different layers in a 2D image generator. Based on this, we divide the T3D face generation process into two parts: NeRF blocks generate coarse textual semantics like facial contour and shape, and 2D upsample blocks render intricate facial attributes. We introduce a Style Code Enhancer (SCE) to enhance fine-grained text semantics incorporation and improve the impact of style codes in the rendering process.
@@
Equipped with Style Code Enhancer, $E^{3}$-FaceNet can better embed fine-grained text information into the rendering process, as it refines and augments missing semantic information based on both the given text and the intermediate visual features while maintaining the image quality.
@@ 4.2 Experimental Results
Comparison with T3D Face methods. We first quantitatively compare our $E^{3}$-FaceNet with two strong T3D Face methods, namely Latent3D (Canfes et al., 2023) and Describe3D (Wu et al., 2023) on CelebAText (Sun et al., 2021) and FFHQ-Text (Zhou, 2021), of which results are given in Tab.1. From this table, we can first find that $E^{3}$-FaceNet greatly outperforms the other two methods under the metric of MVIC, e.g., +6.87% on CelebAText, showing great advantages in terms of multi-view identity preservation. Meanwhile, the performance of SA is also outstanding, which is slightly worse than Latent3D on CelebAText but obviously better on FFHQ. These results suggest that $E^{3}$-FaceNet can achieve excellent multi-view semantic alignment. Moreover, retaining high performance, our $E^{3}$-FaceNet also speeds up the five-view generations by orders of magnitudes. For example, compared with Latent3D and Describe3D, the inference speed of $E^{3}$-FaceNet is 471.75 $\times$ and 35.83 $\times$ faster. Notably, $E^{3}$-FaceNet achieves such performance under the setting of zero-shot generation.
@@
Text-Driven 3D Face Manipulation. To examine the editing ability of $E^{3}$-FaceNet, we also compare it with Latent3D and ClipFace (Aneja et al., 2023) in Tab.2. Without available evaluations, we conduct a user study for this comparison. We can first observe that the identity preservation and the editing quality of $E^{3}$-FaceNet are much superior to the compared methods, e.g., 69.92 v.s. 21.92 of Latent3D on IP and 75.92 v.s. 12.48 of ClipFace on EQ. Besides, $E^{3}$-FaceNet also exhibits strong cross-modal semantic consistency, e.g., 52.64 v.s. 44.56 of ClipFace on SA. In addition, the efficiency of $E^{3}$-FaceNet is still much better than the others. For instance, ClipFace requires about 17 mins to train the texture mappers with another 1.8 seconds for rendering. In contrast, $E^{3}$-FaceNet only takes 0.89 seconds for the five-view editing.
@@ 5 Conclusion
Acknowledgements
This work was supported by National Key R&D Program of China (No.2022ZD0118201), the National Science Fund for Distinguished Young Scholars (No.62025603), the National Natural Science Foundation of China (No. U21B2037, No. U22B2051, No.U21A20472, No. 62072389, No. 623B2088, No. 62176222, No. 62176223, No. 62176226, No. 62072386, No. 62072387, No. 62002305 and No. 62272401), the National Natural Science Fund for Young Scholars of China (No. 62302411), China Postdoctoral Science Foundation (No. 2023M732948), the Natural Science Foundation of Fujian Province of China (No.2021J01002, No.2022J06001), and partially sponsored by CCF-NetEase ThunderFire Innovation Research Funding (NO. CCF-Netease 202301).
@@ Appendix A Introduction of Comparison Methods
TediGAN (Xia et al., 2021) can perform text-guided 2D image generation by optimizing a random sampled noise $z*$ via
@@ B.4.3 Evaluation Criteria
For each editing instruction, we render manipulated 3D faces from three different views: left, front, and right. Participants are presented with the original face and all editing results. We invite 25 individuals to assess the edited faces based on three criteria: (1E) Identity Preservation (IP), selecting the edit result that best preserves the facial identity of the original face; (2) Semantic Alignment (SA), selecting the edit result that most closely aligns with the given instructions; and (3) Editing Quality (EQ), selecting the edit result with the highest perceptual quality. We calculated the percentage of times each method was chosen as the best performer by users, based on the previous mentioned criteria. As shown in Figure 8, the edit results are presented anonymously to the users.